5. Suppose you borrow 100,000 rupees from a bank that charges 15% interest. Using relevant theories, determine how much you will owe the bank over a period of 5 years.
1
Expert's answer
2020-10-07T16:11:21-0400

Discrete compound interest


"A=P(1+r)^t"

"A=100000(1+0.15)^5=201135.72\ rupees"

Discrete compound interest


"A=Pe^{rt}"

"A=100000\cdot e^{0.15\cdot5}=211700.00\ rupees"